OpenStudy (anonymous):

Find the derivative of P(t)=(2)(e^(-2e)^(2t))). Can someone please show me step by step how to solve this?

OpenStudy (kainui):

\[\LARGE P(t)=2e^{-2(e^{2t})}\] Is that it or is it \[\LARGE P(t)=2e^{(-2e)^{2t}}\] I just want to be sure I've got it right.

geerky42 (geerky42):

We need to know. @lana_lei

OpenStudy (anonymous):

It's the second one.

OpenStudy (kainui):

Alright, try your best to do it. You're going to need to use the chain rule here a few times. Are you sure the negative sign is inside the parenthesis and not just the 2 like this: \[\LARGE P(t)=2e^{-(2e)^{2t}}\]

OpenStudy (anonymous):

There aren't any parenthesis in the question, but I just used them to separate each of the exponents. \[P(t) = 2\times e ^{-2e}^{2t}\]I'm really not sure where to start. Would I start by finding the compositions: g(x)=e^e and f(x) = -2e? But what about the 2t exponent?

OpenStudy (kinggeorge):

If there aren't any parentheses, then the equation would be\[\Large P(t)=2e^{-2(e^{2t})}.\]The best way to differentiate this, would be to simply slowly apply the chain rule a few times. So the first equation you have, is\[f(u(t))=2e^{u(t)}\implies f'(u)=2e^{u(t)}\cdot u'(t)\]Now, replace \(u(t)\) with\[-2e^{2t},\]and \(u'(t)\) with\[-4e^{2t}.\]Simplify, and you should be done.

OpenStudy (anonymous):

Thank you so much, this was taking me forever to do! Your explanation was really easy to understand!

OpenStudy (kinggeorge):

You're welcome.